# Misiurewicz points and subhyperbolicity in unicritical algebraic correspondences

### v1

@@ 1. Introduction
In the early 1980s, Douady and Hubbard, drawing on Thurston’s insights into the topological characterization of post-critically finite rational maps [DH93], gave the first definition of sub-hyperbolic maps in [DH84], providing characterizations based on the behavior of critical orbits. In that work, they also carried out a detailed study of polynomial maps of the form $z^{d}+c$ for degrees $d>1$.
1.1. Main results.
@@
Building on Thurston’s orbifold theory [DH93] and elementary sheaf theory, our main contributions are Theorems A, B, C, and D, presented in the next section, where we extend the notion of sub-hyperbolicity to algebraic correspondences of the form (2.1). This setting naturally generalizes the classical family $f_{c}(z)=z^{d}+c$ to $\mathbf{f}_{c}(z)=z^{r}+c$, where $r>1$ is a rational exponent. Our main result establishes that if $a\in\mathbb{C}$ is a Misiurewicz point (Definition 2.1), then $\mathbf{f}_{c}$ expands a suitable orbifold metric uniformly on a neighborhood of the Julia set, with only finitely many singularities. As a consequence, we obtain that the filled Julia set coincides with the Julia set at any such Misiurewicz point $a$. This more general framework, rooted in the theory of algebraic correspondences, has been the subject of significant attention in recent decades.
1.2. Key developments on algebraic correspondences
@@
The advent of computer graphics in the late 1970s, together with Mandelbrot’s iconic visualizations of fractals, sparked a global interest in holomorphic dynamics. This revival also brought renewed attention to multi-valued systems (long studied in the context of Fuchsian groups), particularly following Bullett’s work in the early 1990s on critically finite correspondences [Bullett92], and the subsequent breakthrough by Bullett and Penrose, who introduced a family $\mathcal{F}_{a}$ of algebraic correspondences (matings) whose connectedness locus was conjectured to be homeomorphic to the Mandelbrot set [Bullett1994].
@@
Nearly three decades later, this conjecture was confirmed through a series of major advances. Bullett and Lomonaco [BL19, BULLETT2024109956] provided a rigorous proof of the conjecture originally posed in [Bullett1994], while independent developments by Lee, Lyubich, Makarov, and Mukherjee [Mukherjee] introduced a family $\mathcal{S}$ of Schwarz reflections that yield anti-holomorphic correspondences, realized as matings between anti-rational maps and the abstract modular group. One of the most striking results in [Mukherjee] is the construction of a homeomorphism between the abstract connectedness locus of $\mathcal{S}$ and the abstract parabolic Tricorn, a combinatorial model of the Tricorn. The latter can be interpreted as the connectedness locus of a family of anti-holomorphic quadratic polynomials, also known as the anti-holomorphic Mandelbrot set.
1.3. Application to asymptotic similarity.
@@
Our results serve as key tools for analyzing the similarity between Multibrot sets and Julia sets arising from the dynamics of $\mathbf{f}_{c}$. Just as Tan Lei [TanLei] established an analogous phenomenon for the quadratic family $z^{2}+c$, building on earlier results developed by Douady and Hubbard [DH84], we will apply the theorems of this paper to extend Tan Lei’s results to correspondences $\mathbf{f}_{c}(z)=z^{r}+c$, where $r>1$ is rational. This generalization is carried out in detail in [siqueira2025quadratic]. See also Figure 1 for an illustration in this context.
1.4. On critically finite correspondences.
@@
It is worth noting that the application of Thurston’s classification to the setting of algebraic correspondences was first explored in the seminal work of Bullett [Bullett92], who classified critically finite quadratic correspondences (those in which every critical point has a finite full orbit under both forward and backward iteration) showing that such systems exhibit strong rigidity. In particular, Bullett demonstrated that, up to conformal conjugacy, there exist only eleven such correspondences of quadratic type.
@@ 2. Definitions and main results
Recall that for the quadratic family $f_{c}(z)=z^{2}+c$, a point $c\in\mathbb{C}$ is called a Misiurewicz point if the critical point is strictly pre-periodic under iterations by $f_{c}.$ Thanks to special geometric properties of $f_{c}$, it can be shown that the critical point eventually maps to a repelling cycle when the parameter $c$ is a Misiurewicz point. As previously described in [ETDS22, SS17, Proc22], many of the geometric features characteristic of the quadratic family – such as rigidity, hyperbolic Julia sets of zero area, and holomorphic motions – also extend to maps of the form $z^{r}+c,$ where $r>1$ is a rational number. However, such maps are no longer single-valued, but rather define algebraic correspondences given by
@@
and $(ii)$ this orbit is strictly pre-periodic. The first point of this orbit which is periodic is denoted by $z_{\ell}$ and the associated cycle $\alpha_{a}$ of period $n$ is
@@
Repelling cycles are defined in section $2.$ Theorem D shows that the cycle $\alpha_{a}$ is, in fact, repelling.
@@
Let $p,q$ be positive integers with $p\geq q>1$. The Multibrot set $M_{p,q}$ associated with the family (2.1) is defined by
@@
For the quadratic family $f_{c}(z)=z^{2}+c$, Misiurewicz points form a countable and dense subset of the boundary of the Mandelbrot set; see [milnor1989self].
@@
We will prove that there exist infinitely many Misiurewicz points in a neighborhood of c=2c=-2 for the semigroup family $\langle z^{2}+c,\ -z^{2}+c\rangle$. The parameter c=2c=-2 is a Misiurewicz point for this family: we will show that the critical point 0 has only one bounded forward orbit, which is strictly pre-periodic, mapping to the fixed point $2.$
@@
Figure 1. The Multibrot set $M_{4,2}$ and related copies, illustrating the similarity between $M_{4,2}$ and the filled Julia set $K_{-2}$ near the parameter 2-2. There exists an infinite sequence of Misiurewicz points $c_{n}$, each located near a small copy of $M_{4,2}$, and converging to 2-2. See Example 2.1. The point 2-2 is not exceptional; similar patterns occur near infinitely many copies of $M_{4,2}$, as well as in other Multibrot sets $M_{p,q}$, as discussed in [siqueira2025quadratic].
@@
On the other hand, Misiurewicz parameters are dense in the boundary $\partial M$ of the Mandelbrot set for the quadratic family $z^{2}+c$, so there exists a sequence $c_{n}\to-2$ for which the orbit of 0 under $z^{2}+c_{n}$ is strictly pre-periodic. Since this orbit is also admissible under the correspondence, each $c_{n}$ admits exactly one bounded, strictly pre-periodic orbit of the critical point under $\mathbf{f}_{c_{n}}$. Hence each $c_{n}$ is a Misiurewicz point for the correspondence.
In contrast, although $c=i$ is a Misiurewicz parameter for the quadratic family, it fails to satisfy the same property for the correspondence, as the critical point has multiple bounded forward orbits and is not pre-periodic to a single repelling cycle. Indeed, after computing $\mathbf{f}_{i}^{3}(i)$, we find that only four orbits of 0 remain bounded, while all others escape to the basin of infinity. These four distinct orbits eventually land on four different cycles, so the critical point is not pre-periodic to a single cycle. Thus $i$ is not a Misiurewicz parameter for the correspondence.
@@
It should be noted that, according to the results in [ETDS22, §2.1], the filled Julia set $K_{c}$ is connected for every parameter $c\in M_{p,q}$. The set $M_{2,1}$ coincides with the classical Mandelbrot set.
The following theory, due to Thurston [DH93], is both elegant and deserving of a clear presentation, as it underpins our main results.
2.1. Branched coverings and orbifolds
An orbifold is a pair $(X,\nu)$ where $X$ is a Riemann surface and $\nu:S\to\{1,2,\ldots\}$ is a function such that the set of ramified points $\{x\in X:\nu(x)>1\}$ is locally finite. The integer $\nu(x)$ is the branch index of $x.$
Suppose that $X$ and $Y$ are Riemann surfaces. A surjective holomorphic map $\varphi:X\to Y$ is proper if the inverse image of any compact set in $Y$ is a compact set in $X.$ Proper maps satisfy some interesting properties: the set of branch points $B=\{x\in X:\varphi^{\prime}(x)=0\}$ is locally finite, as well as $R=\varphi(B)$ and $\varphi^{-1}(R)$. The induced map $f:X{\setminus}\varphi^{-1}(R)\to X{\setminus}R$ is a covering map of finite degree $d$. For this reason proper maps are also known as d-fold branched coverings.
@@ 2.1. Branched coverings and orbifolds
A few properties of regular branched coverings: the set of branch point $B$ is locally finite and $\Gamma.B=B.$ The local degree of $\varphi$ at a point $x\in B$ is $n$ if $\varphi^{(k)}(x)=0$ for all $1\leq k<n$, and $\varphi^{(n)}(x)\neq 0$. The local degree of $\varphi$ at $x\in B$ is the same as the local degree of $\varphi$ at $\gamma.x,$ for all $\gamma\in\Gamma.$ Any point of $R=\varphi(B)$ is a ramified point. Since $\varphi$ is regular, $\varphi^{-1}(R)=B;$ moreover, the local degree of $\varphi$ at every point in the inverse image of a ramified point $y$ is the same integer $d$, defined as the ramification index $\nu(y)=d.$ This defines a ramification function $\nu:Y\to\{1,2,\ldots\}$ associated with the regular branched covering $\varphi$. Notice that $y$ is a ramified point if, and only if, $\nu(y)>1.$
@@
Let $(S,\nu)$ be a Riemann surface orbifold which is conformally isomorphic to the complex plane. Suppose that $\nu$ is non-trivial, that is, we have at least two ramified points with different ramification indexes. Then there exists a regular branched covering $\varphi:\tilde{S}_{\nu}\to(\mathbb{C},\nu)$, unique up to conformal isomorphisms, such that $\tilde{S}_{\nu}$ is conformally isomorphic to the hyperbolic disk $\mathbb{D}$ and the ramification function of this covering is the given $\nu.$
@@
Our statement follows the exposition of John Milnor in [Milnor], which itself is based on the original theory developed in the seminal paper [DH93]. The result discussed here is a special case of Theorem E.1 in [Milnor], with the additional conclusion that the universal covering space is $\tilde{S}_{\nu}\simeq\mathbb{D}$, a fact we now proceed to justify.
According to Douady and Hubbard [DH93], the Euler characteristic of the orbifold $(S,\nu)$ is defined by
@@
It then follows from Lemmas E.3 and E.4 in [Milnor] that the universal cover $\tilde{S}_{\nu}$ is hyperbolic; that is, $\tilde{S}_{\nu}\simeq\mathbb{D}$. ∎
Let $(\mathbb{C},\nu)$ be an orbifold with a non-trivial ramification function $\nu.$ Let $\varphi$ denote the regular branched covering given by Theorem 2.1. As usual, the set of branch points is $B=\{x\in\tilde{S}_{\nu}:\varphi^{\prime}(x)=0\}$ and its image under $\varphi$ is the set of ramified points, which we denote by $R.$ Since $\varphi$ is regular, $\varphi^{-1}(R)=B.$ Both $B$ and $R$ are locally finite. The restriction $\varphi:\tilde{S}_{\nu}{\setminus}B\to\mathbb{C}{\setminus}R$ is a covering map. Let $\rho_{1}(z)|dz|$ denote the Poincaré metric of $\tilde{S}_{\nu}\simeq\mathbb{D}.$ Every element of the associated group $\Gamma$ of deck transformations is an isometry with respect to $\rho_{1}(z)|dz|.$ Using this fact we are allowed to make the following definition.
@@
There exists a unique conformal metric $\rho_{2}(z)|dz|$ on $\mathbb{C}{\setminus}R$ for which
@@
The orbifold metric blows up at each ramified point $a_{j}$: in some punctured neighborhood $U^{*}$ of $a_{j},$ we have $\rho_{2}(z)\to\infty$ as $z\to a_{j}.$ (This property can be checked on page 211 of [Milnor]. Note that the local branched covering $z=a_{j}+w^{\nu(a_{j})}$ and $\varphi$ have the same ramification index at $a_{j}$).
@@
where $\mathbf{f}_{a}^{0}$ is the identity. Since this set is countable, we will denote its elements by $a_{j}$, with $a_{0}=0$ and $j\geq 0.$
@@
The ramification function $\nu_{a}$ associated to the Misiurewicz point $a$ is defined by setting $\nu_{a}(0)=q,$ $\nu_{a}(a_{j})=p$ for every nonzero point $a_{j}$ in the post-critical set, and $\nu_{a}(z)=1$ elsewhere. It is always assumed that $q\geq 2$ and $p>q.$ The pair $(\mathbb{C},\nu_{a})$ is the canonical orbifold of the correspondence. The canonical orbifold metric of $(\mathbb{C},\nu_{a})$ is the one given in Definition 2.4 with $\nu=\nu_{a}.$
The following theorem is central to this work. To make its terminology precise, we recall in Section 3 key notions from global analytic functions.
@@
Proof.
See section 4. ∎
Constructing global branches of the correspondence on the complex plane is, in general, obstructed by the presence of the algebraic singularity at zero. Remarkably, after lifting $\mathbf{f}_{c}$ to the unit disk, such branches can indeed be defined, as the following theorem shows.
@@
Suppose that $a$ is a Misiurewicz point of the family $\mathbf{f}_{c}$ given by (2.1). Let $\mathbf{g}$ be the holomorphic multifunction of Theorem A, which is defined on the open unit disk. Every branch $f$ of $\mathbf{g}$ has a unique extension to a global branch $F:\mathbb{D}\to\mathbb{D}.$ Moreover, $F$ strictly contracts the Poincaré metric. Since $\varphi\circ F$ is a branch of $\mathbf{f}_{a}^{-1}\circ\varphi,$ the following diagram is commutative:
@@
We know from [ETDS22, equation $(3)$] that the filled Julia set $K_{c}$ is compact. Since all repelling cycles lie within $K_{c}$, it follows that $J_{c}\subset K_{c}$.
@@
Our proof that $J_{a}=K_{a}$ in this setting relies on the sub-hyperbolicity of $\mathbf{f}_{a}$ and the construction of homoclinic orbits. Alternative approaches, obtained by straightforward extensions of the standard arguments used to establish $J_{a}=K_{a}$ for polynomials, fail in the case of correspondences.
3. Background on global analytic functions
The content of this section summarizes key results from the classical exposition in L. Ahlfors [Ahlfors, Chapter 8]. While it is primarily based on that source, our presentation has the advantage of being more concise, allowing the reader to grasp the essential notation and ideas in just a few pages. Note that the notation used here differs slightly from that of Ahlfors. In addition, we have introduced new definitions and related topics that do not appear in Ahlfors (or, to our knowledge, in any other standard exposition of the subject). For these results, we provide brief proofs.

### v2

@@ 1. Introduction
In the early 1980s, Douady and Hubbard, drawing on Thurston’s insights into the topological characterization of postcritically finite rational maps [7], gave the first definition of subhyperbolic maps in [8], providing characterizations based on the behavior of critical orbits. In that work, they also carried out a detailed study of polynomial maps of the form $z^{d}+c$ for degrees $d>1$.
Main results.
@@
Building on Thurston’s orbifold theory [7] and elementary sheaf theory, our main contributions are Theorems A, B, C, and D, presented in the next section, where we extend the notion of subhyperbolicity to algebraic correspondences of the form (2.1). This setting naturally generalizes the classical family $f_{c}(z)=z^{d}+c$ to $\mathbf{f}_{c}(z)=z^{r}+c$, where $r>1$ is a rational exponent. Our main result establishes that if $a\in\mathbb{C}$ is a Misiurewicz point (Definition 2.1), then $\mathbf{f}_{c}$ expands a suitable orbifold metric uniformly on a neighborhood of the Julia set, with only finitely many singularities. As a consequence, we obtain that the filled Julia set coincides with the Julia set at any such Misiurewicz point $a$. This more general framework, rooted in the theory of algebraic correspondences, has been the subject of significant attention in recent decades.
Key developments on algebraic correspondences
@@
The advent of computer graphics in the late 1970s, together with Mandelbrot’s iconic visualizations of fractals, sparked a global interest in holomorphic dynamics. This revival also brought renewed attention to multi-valued systems (long studied in the context of Fuchsian groups), particularly following Bullett’s work in the early 1990s on critically finite correspondences [6], and the subsequent breakthrough by Bullett and Penrose, who introduced a family $\mathcal{F}_{a}$ of algebraic correspondences (matings) whose connectedness locus was conjectured to be homeomorphic to the Mandelbrot set [5].
@@
Nearly three decades later, this conjecture was confirmed through a series of major advances. Bullett and Lomonaco [2, 4] provided a rigorous proof of the conjecture originally posed in [5], while independent developments by Lee, Lyubich, Makarov, and Mukherjee [9] introduced a family $\mathcal{S}$ of Schwarz reflections that yield anti-holomorphic correspondences, realized as matings between anti-rational maps and the abstract modular group. One of the most striking results in [9] is the construction of a homeomorphism between the abstract connectedness locus of $\mathcal{S}$ and the abstract parabolic Tricorn, the combinatorial model of the Tricorn. The latter can be interpreted as the connectedness locus of a family of anti-holomorphic quadratic polynomials, also known as the anti-holomorphic Mandelbrot set.
Application to asymptotic similarity.
@@
Our results serve as key tools for analyzing the similarity between Multibrot sets and Julia sets arising from the dynamics of $\mathbf{f}_{c}$. Just as Tan Lei [10] established an analogous phenomenon for the quadratic family $z^{2}+c$, building on earlier results developed by Douady and Hubbard [8], we will apply the theorems of this paper to extend Tan Lei’s results to correspondences $\mathbf{f}_{c}(z)=z^{r}+c$, where $r>1$ is rational. This generalization is carried out in detail in [16]. See also Figure 1 for an illustration in this context.
On critically finite correspondences.
@@
It is worth noting that the application of Thurston’s classification to the setting of algebraic correspondences was first explored in the seminal work of Bullett [6], who classified critically finite quadratic correspondences (those in which every critical point has a finite full orbit under both forward and backward iteration), showing that such systems exhibit strong rigidity. In particular, Bullett demonstrated that, up to conformal conjugacy, there exist only eleven such correspondences of quadratic type.
@@ 2. Definitions and main results
Recall that for the quadratic family $f_{c}(z)=z^{2}+c$, a point $c\in\mathbb{C}$ is called Misiurewicz if the critical point is strictly preperiodic under iterations by $f_{c}.$ Thanks to special geometric properties of $f_{c}$, it can be shown that the critical point eventually maps to a repelling cycle when the parameter $c$ is a Misiurewicz point. As previously described in [14, 13, 15], many of the geometric features characteristic of the quadratic family – such as rigidity, hyperbolic Julia sets of zero area, and holomorphic motions – also extend to maps of the form $z^{r}+c,$ where $r=p/q>1$ is a rational number. However, such maps are no longer single-valued, but rather define algebraic correspondences given by
@@
and $(ii)$ this orbit is strictly preperiodic. The first point of this orbit which is periodic is denoted by $z_{\ell}$ and the associated cycle $\alpha_{a}$ of period $n$ is
@@
Repelling cycles are defined in Section 2. Theorem D shows that the cycle $\alpha_{a}$ is repelling.
The motivation for the following definition is discussed in Remark 2.4.
@@
Let $p,q$ be positive integers with $p>q\geq 1$. The Multibrot set $M_{p,q}$ associated with the family (2.1) is defined by
@@
For the quadratic family $f_{c}(z)=z^{2}+c$, Misiurewicz points form a countable and dense subset of the boundary of the Mandelbrot set; see [11].
@@
We will prove that there exist infinitely many Misiurewicz points in a neighborhood of c=2c=-2 for the semigroup family $\langle z^{2}+c,\ -z^{2}+c\rangle$. The parameter c=2c=-2 is a Misiurewicz point for this family: we will show that the critical point 0 has only one bounded forward orbit, which is strictly preperiodic, mapping to the fixed point $2.$
@@
Figure 1. The Multibrot set $M_{4,2}$ and related copies, illustrating the similarity between $M_{4,2}$ and the filled Julia set $K_{-2}$ near the parameter 2-2. There exists an infinite sequence of Misiurewicz points $c_{n}$, each located near a small copy of $M_{4,2}$, and converging to 2-2. See Example 2.1. The point 2-2 is not exceptional; similar patterns occur near infinitely many copies of $M_{4,2}$, as well as in other Multibrot sets $M_{p,q}$, as discussed in [16].
@@
On the other hand, Misiurewicz parameters are dense in the boundary $\partial M$ of the Mandelbrot set for the quadratic family $z^{2}+c$, so there exists a sequence $c_{n}\to-2$ for which the orbit of 0 under $z^{2}+c_{n}$ is strictly preperiodic. Since this orbit is also admissible under the correspondence, each $c_{n}$ admits exactly one bounded, strictly preperiodic orbit of the critical point under $\mathbf{f}_{c_{n}}$. Hence each $c_{n}$ is a Misiurewicz point for the correspondence.
In contrast, although $c=i$ is a Misiurewicz parameter for the quadratic family, it fails to satisfy the same property for the correspondence, as the critical point has multiple bounded forward orbits and is not preperiodic to a single repelling cycle. Indeed, after computing the iterate $\mathbf{f}_{i}^{3}(i)$, we find that only four orbits of 0 remain bounded, while all others escape to the basin of infinity. These four distinct orbits eventually land on four different cycles, so the critical point is not preperiodic to a single cycle. Thus $i$ is not a Misiurewicz parameter for the correspondence.
@@
It should be noted that, according to the results in [14, §2.1], the filled Julia set $K_{c}$ is connected for every parameter $c\in M_{p,q}$. The set $M_{2,1}$ coincides with the classical Mandelbrot set.
Remark 2.2.
Consider the quadratic family $f_{c}(z)=z^{2}+c$. If $a\in\mathbb{C}$ is a Misiurewicz parameter, it is well known that:
  1. (1)
    $J_{a}=K_{a}$, and the filled Julia set $K_{a}$ has empty interior;
  2. (2)
    there exist infinitely many small copies of the Mandelbrot set accumulating at $a$;
  3. (3)
    the Mandelbrot set $M$ and the filled Julia set $K_{a}$ are asymptotically similar in a neighborhood of $a$.
As shown in Example 2.1, when $c=i$ the critical point is preperiodic to multiple distinct cycles under the dynamics of the semigroup $\langle z^{2}+c,\,-z^{2}+c\rangle$. In this situation, experiments indicate that the associated filled Julia set satisfies none of the properties (1)(1)$(3).$ These observations motivate our definition of a Misiurewicz point. Allowing the critical point to be preperiodic to multiple cycles may lead to interesting phenomena; however, such behavior does not generalize the classical polynomial case.
The following theory, due to Thurston [7], is both elegant and deserving of a clear presentation, as it underpins our main results.
Branched coverings and orbifolds
An orbifold is a pair $(X,\nu)$ where $X$ is a Riemann surface and $\nu:X\to\{1,2,\ldots\}$ is a function such that the set of ramified points $\{x\in X:\nu(x)>1\}$ is locally finite. The integer $\nu(x)$ is the branch index of $x.$
Suppose that $X$ and $Y$ are Riemann surfaces. A surjective holomorphic map $\varphi:X\to Y$ is proper if the inverse image of any compact set in $Y$ is a compact set in $X.$ Proper maps satisfy some interesting properties: the set of branch points $B=\{x\in X:\varphi^{\prime}(x)=0\}$ is locally finite, as well as $R=\varphi(B)$ and $\varphi^{-1}(R)$. The induced map $f:X{\setminus}\varphi^{-1}(R)\to Y{\setminus}R$ is a covering map of finite degree $d$. For this reason proper maps are also known as $d$-fold branched coverings.
@@
A few properties of regular branched coverings: the set of branch points $B$ is locally finite and $\Gamma.B=B.$ The local degree of $\varphi$ at a point $x\in B$ is $n$ if $\varphi^{(k)}(x)=0$ for all $1\leq k<n$, and $\varphi^{(n)}(x)\neq 0$. The local degree of $\varphi$ at $x\in B$ is the same as the local degree of $\varphi$ at $\gamma.x,$ for all $\gamma\in\Gamma.$ Any point of $R=\varphi(B)$ is a ramified point. Since $\varphi$ is regular, $\varphi^{-1}(R)=B;$ moreover, the local degree of $\varphi$ at every point in the inverse image of a ramified point $y$ is the same integer $d$, defined as the ramification index $\nu(y)=d.$ This defines a ramification function $\nu:Y\to\{1,2,\ldots\}$ associated with the regular branched covering $\varphi$. Notice that $y$ is a ramified point if, and only if, $\nu(y)>1.$
@@
Let $(S,\nu)$ be a Riemann surface orbifold that is conformally isomorphic to the complex plane. Suppose that $\nu$ is non-trivial, that is, we have at least two ramified points with different ramification indices. Then there exists a regular branched covering $\varphi:\tilde{S}_{\nu}\to(\mathbb{C},\nu)$, unique up to conformal isomorphisms, such that $\tilde{S}_{\nu}$ is conformally isomorphic to the hyperbolic disk $\mathbb{D}$ and the ramification function of this covering is the given $\nu.$
@@
Our statement follows the exposition of John Milnor in [12], which itself is based on the original theory developed in the seminal paper [7]. The result discussed here is a special case of Theorem E.1 in [12], with the additional conclusion that the universal covering space is $\tilde{S}_{\nu}\simeq\mathbb{D}$, a fact we now proceed to justify.
According to Douady and Hubbard [7], the Euler characteristic of the orbifold $(S,\nu)$ is defined by
@@
It then follows from Lemmas E.3 and E.4 in [12] that the universal cover $\tilde{S}_{\nu}$ is hyperbolic; that is, $\tilde{S}_{\nu}\simeq\mathbb{D}$. ∎
Let $(\mathbb{C},\nu)$ be an orbifold with a non-trivial ramification function. Let $\varphi$ denote the regular branched covering given by Theorem 2.1.
Let $\varphi$ denote the regular branched covering given by Theorem 2.1. As usual, the set of branch points is $B=\{x\in\tilde{S}_{\nu}:\varphi^{\prime}(x)=0\}$ and its image under $\varphi$ is the set of ramified points, which we denote by $R.$ Since $\varphi$ is regular, $\varphi^{-1}(R)=B.$ Both $B$ and $R$ are locally finite. The restriction $\varphi:\tilde{S}_{\nu}{\setminus}B\to\mathbb{C}{\setminus}R$ is a covering map. Let $\rho_{1}(z)|dz|$ denote the Poincaré metric on $\tilde{S}_{\nu}\simeq\mathbb{D}$. Every element of the associated group $\Gamma$ of deck transformations is an isometry with respect to $\rho_{1}(z)|dz|.$ Using this fact we are allowed to make the following definition.
@@
There exists a unique conformal metric $\rho_{2}(z)|dz|$ on $\mathbb{C}{\setminus}R$ such that
@@
The orbifold metric blows up at each ramified point $a_{j}$: in some punctured neighborhood $U^{*}$ of $a_{j},$ we have $\rho_{2}(z)\to\infty$ as $z\to a_{j}.$ (This property can be verified on page 211 of [12]. Note that the local branched covering $z=a_{j}+w^{\nu(a_{j})}$ and $\varphi$ have the same ramification index at $a_{j}$.)
@@
where $\mathbf{f}_{a}^{0}$ denotes the identity map. Since this set is countable, we will denote its elements by $a_{j}$, with $a_{0}=0$ and $j\geq 0.$
@@
The ramification function $\nu_{a}$ associated with the Misiurewicz point $a$ is defined by setting $\nu_{a}(0)=q,$ $\nu_{a}(a_{j})=p$ for every nonzero point $a_{j}$ in the postcritical set, and $\nu_{a}(z)=1$ elsewhere. It is always assumed that $q\geq 2$ and $p>q.$ The pair $(\mathbb{C},\nu_{a})$ is the canonical orbifold of the correspondence. The canonical orbifold metric of $(\mathbb{C},\nu_{a})$ is the one given in Definition 2.4 with $\nu=\nu_{a}.$
The sheaf $\mathfrak{S}_{U}$
The following theorem is central to this work. In order to state it precisely, we briefly recall some standard notions from the theory of global analytic functions. Further background and related discussion can be found in Section 3.
Let $U\subset\mathbb{C}$ be a domain. The set of all germs $(f,z)$ of holomorphic functions with base point $z\in U$, denoted by $\mathfrak{S}_{U}$, carries a natural topology, under which it becomes the sheaf of germs of analytic functions on $U$. Each connected component of this space is a Riemann surface, and the canonical projection $\pi\colon(f,z)\mapsto z$ is a local homeomorphism.
Given a germ $(f,z)$ and a continuous path $\gamma\colon[0,1]\to U$ with $\gamma(0)=z$, an analytic continuation of $(f,z)$ along $\gamma$ is a continuous lift $\widetilde{\gamma}\colon[0,1]\to\mathfrak{S}_{U}$ such that $\widetilde{\gamma}(0)=(f,z)$ and
$\pi\circ\widetilde{\gamma}(t)=\gamma(t)\quad\text{for all }t\in[0,1].$
A multifunction $\mathbf{f}$ is a set–valued mapping defined on a domain $U\subset\mathbb{C}$, assigning to each point $z\in U$ a subset $\mathbf{f}(z)\subset\mathbb{C}$. A branch of $\mathbf{f}$ is a holomorphic map $f$ such that $f(z)\in\mathbf{f}(z)$ for every $z$ in the domain of $f$. We say that $\mathbf{f}$ is a holomorphic multifunction if, for every $z\in U$ and every $w\in\mathbf{f}(z)$, there exists a branch $f$ of $\mathbf{f}$ locally defined at $z$ such that $f(z)=w$.
Given a holomorphic multifunction $\mathbf{f}\colon U\to\mathbb{C}$, we denote by $\mathfrak{S}(\mathbf{f})$ the space of all germs $\mathfrak{f}=(f,z)$ arising from local branches $f$ of $\mathbf{f}$. If $\mathfrak{S}(\mathbf{f})$ is a connected component of the sheaf of germs of analytic functions on $U$, then $\mathbf{f}$ is called a global analytic multifunction.
@@
Constructing global branches of the correspondence on the complex plane is, in general, obstructed by the presence of the algebraic singularity at zero. Remarkably, after lifting $\mathbf{f}_{c}^{-1}$ to the unit disk, such branches can indeed be defined, as the following theorem shows.
@@
Suppose that $a$ is a Misiurewicz point of the family $\mathbf{f}_{c}$ given by (2.1). Let $\mathbf{g}$ be the holomorphic multifunction of Theorem A, defined on the open unit disk. Every branch $f$ of $\mathbf{g}$ has a unique extension to a global branch $F\colon\mathbb{D}\to\mathbb{D}$. Moreover, $F$ is strictly contracting with respect to the Poincaré metric. Since $\varphi\circ F$ is a branch of $\mathbf{f}_{a}^{-1}\circ\varphi$, the following diagram commutes:
@@
Remark 2.4.
By [14, equation (3)], the filled Julia set $K_{c}$ is given by the nested intersection
$K_{c}=\bigcap_{k\geq 0}\mathbf{f}_{c}^{-k}(D),$
where $D=\{\,|z|\leq R\,\}$ is any sufficiently large disk. Thus $K_{c}$ is defined in complete analogy with the filled Julia set of a polynomial. An immediate consequence is that $K_{c}$ is compact and contains all cycles, including repelling ones. Hence the Julia set $J_{c}$, defined as the closure of the repelling cycles, satisfies $J_{c}\subset K_{c}$.
For polynomials one has $J_{c}=\partial K_{c}$. In contrast, except in the trivial case $c=0$ (where $J_{0}$ is the unit circle and $K_{0}$ the unit disk) in general $\partial K_{c}$ is a proper subset of $J_{c}$, with parts of $J_{c}$ lying in the interior of $K_{c}$. See [3, Section 3] for a more detailed discussion of the geometry of $K_{c}$ and $J_{c}$.
@@
Our proof that $J_{a}=K_{a}$ in this setting relies on the subhyperbolicity of $\mathbf{f}_{a}$ and the construction of homoclinic orbits. Alternative approaches, obtained by straightforward extensions of the standard arguments used to establish $J_{a}=K_{a}$ for polynomials, fail in the case of correspondences.
Outline.
The structure of this paper is as follows. Sections 1 and 2 present the main results and related topics. Section 3 discusses analytic continuation; while it contains no new results, it should not be skipped, as its primary purpose is to establish precise and consistent notation. The proof of Theorem A is given in Section 4, and Theorem B is proved in Section 5. Section 6 establishes uniform contraction of branches of $\mathbf{f}_{a}^{-1}$ on compact sets, from which Theorem C follows; its proof appears in Section 7. Section 7 also contains the proof of Theorem D, the final result of this paper.
3. The sheaf of germs of analytic functions
The content of this section summarizes key results from the classical exposition in L. Ahlfors [1, Chapter 8]. While it is primarily based on that source, our presentation has the advantage of being more concise, allowing the reader to grasp the essential notation and ideas in just a few pages. Note that the notation used here differs slightly from that of Ahlfors. In addition, we have introduced new definitions and related topics that do not appear in Ahlfors (or, to our knowledge, in any other standard exposition of the subject). For these results, we provide brief proofs.